OpenStudy (anonymous):

The length of a rectangle is 5 more than the width. The perimeter of the rectangle is 168 inches. Find the dimensions of the rectangle.

OpenStudy (anonymous):

perimeter = 2l + 2w = 2(w+5) + 2w 2w + 10 + 2w = 168 4w = 158 w = 39.5 width = 39.5 and length = 44.5